A STUDENT WAS ASKED TO DIVIDE 7.5 BY 8.3 AND TO OBTAIN THE QUOTIENT CORRECT TO THREE DECIMAL PLACES, BUT HE DIVIDED 8.3 BY 7.5 AND GOT THE RESULT CORRECT 2 DECIMAL PLACES. BY HOW MUCH PERCENT WAS HIS RESULT WRONG?
PLEASE PROVIDE THE SOLUTION....ANSWER IS 22.8%...

Answer by rothauserc(4718)

7.5 / 8.3 = 0.904
8.3 / 7.5 = 1.11
1.11 - 0.904 = 0.206
0.206 / 0.904 = 0.227876106 approx 22.8%